How do you slove for x on this problem, 2x + 15 = 27 - 4x?
Answer by shankystud19(2)

2x +15 =27-4x
by placing cofficent of x in same side and others terms on other side we get
2x+4x=27-15
6x=12
x=12/6
x=2